# Measuring kinetic inductance and superfluid stiffness of two-dimensional superconductors using high-quality transmission-line resonators

### v2

@@ I Introduction
Precise measurement of complex conductivity can act as a powerful tool to study the fundamental properties of superconductorsTinkham (1996). The reactive component of the conductivity is the kinetic inductance, which arises from the kinetic energy of Cooper pairs. Kinetic inductance can directly measure superfluid stiffness, and therefore, has the capability to reveal quantum geometric contributions to stiffness which are invisible at the level of band dispersion Tovmasyan et al. (2016); Julku et al. (2016); Verma et al. (2021); Törmä et al. (2022), and for which there only exists indirect evidence at presentTian et al. (2023) in twisted graphene. Sensitive measurements of the superfluid stiffness can also be used to detect the weak supercurrent dependence of the penetration depth (known as the nonlinear Meissner effectYip and Sauls (1992); Xu et al. (1995); Stojković and Valls (1995); Sauls (2022)), which offers highly constraining insights into the superconducting pairing symmetry. On the other hand, the resistive component of the complex conductivity reflects the energy dissipation due to normal-state carriers under the framework of the two-fluid model. Finally, kinetic inductance is an important ingredient in superconducting quantum circuitry: it can be used to engineer qubitsWinkel et al. (2020), to realize high-sensitivity photon detectorsDay et al. (2003), and to parametrically amplify signals with performance constrained solely by quantum noiseEom et al. (2012); Ranzani et al. (2018).
@@
In this paper, we describe a new technique for accurately determining the kinetic inductance of these thin, small superconductor sample components. Inspired by the use of superconducting resonators in the readout of qubit states in circuit quantum electrodynamics (cQED), our method is a mesoscopic hybrid setup in which the kinetic inductance of our sample component acts as a small perturbation to the probe resonator. We then extract the relevant sample component parameters from the change of the resonator frequency and quality factor. This contrasts with previous kinetic inductance measurements in which a 50 $\Omega$ resistor shunts out the resonator in order to impedance match to the measurement circuitry, thus reducing the quality ($Q$) factor and sacrificing measurement sensitivityAnnunziata et al. (2010); Singh et al. (2018). Our method shares the same fundamental mechanism as kinetic inductance detectors, which measure the suppression of resonance frequency due to the breaking of Cooper pairs in a superconducting resonator Visser et al. (2014); Day et al. (2003). Already, this method has been applied to study the superfluid stiffness in superconductor-ferromagnet bilayersBøttcher et al. (2023) and FeSe1-xTexCherpak et al. (2023), as well as metrological studies of the loss tangent of materials and the cleanliness of fabrication procedures to advance the development of superconducting quantum technologySage et al. (2011); Quintana et al. (2014); Grünhaupt et al. (2018); McRae et al. (2021); Antony et al. (2022).
@@ II Theory of Kinetic Inductance and Superfluid Stiffness
In a superconductor, kinetic inductance arises from the inertial response of Cooper pairs to an applied current. If all the Cooper pairs contribute to the supercurrent, we can equate the total kinetic energy to kinetic inductance and derive
@@
Apart from spatial profile of ${\bf A}({\bf r})$, we have also assumed the locality of the Mattis-Bardeen kernel in deriving the expression above, i.e. that the current density is proportional to the local vector potential, that is, $J({\bf r})\propto\mathcal{D}_{s}{\bf A}({\bf r})$. In a superconductor with a large coherence length, the electromagnetic response can deviate from the local value of the vector potential. The current spreads over a region of size determined by the coherence length $\xi$ and the Mattis-Bardeen kernel $K({\bf r})$, that is,
[…]
and the resulting energy stored in the superconductor takes the form
[…]
thus modifying the free energy in Eqn. 2. These modifications not only complicate the analysis, they obscure the connection between kinetic inductance, penetration depth and superfluid stiffness. Specifically, superfluid stiffness is defined as the ${\bf q}=0$ component of the Fourier transform $K({\bf q})$ Tinkham (1996), penetration depth is defined from the solution of the non-local London equation Tinkham (1996), and kinetic inductance measures an effective penetration depth that depends on $K({\bf q})$ as well as the boundary conditions Gao (2008).
It is instructive to analyze the locality assumption in Fourier space. The kernel $K({\bf q})$ is constant in ${\bf q}$ space if the response is local. Non-locality brings in a variation at the scale of coherence length $\xi$. Therefore, if $q\xi\ll 1$, we can safely assume a local response. As we outline in Appendix VII.1, sample thickness and the electronic mean free path provide a natural cut-off that works in favor of local response. One may heuristically think of there being an effective coherence length, which is dominated by the smallest of the relevant length scales. As a result, despite the complications, the local approximation has widespread applicability, especially in thin samples.
@@ III.1 Quarter-wave resonator circuit model
where $k$ is the wavenumber. The approximation in Eqn. 7 uses $|r+jx|\ll Z_{0}$, so that the sample component introduces only a small perturbation to the quarter-wave resonator impedance. By comparing the transmission line resonator with the parallel RLC resonator and relating their physical parameters (see Appendix VII.2), the loss becomes $\alpha\ell+r/Z_{0}$ with a frequency shift, $\delta\omega_{0}$, given by:
@@
The resonant-frequency shift is negative for an inductive sample, which effectively elongates the electrical length of the resonator.
@@
This $Q_{\rm eff}$ term has two contributions: the loss tangent of the sample component $\tan{\delta}$, and the participation ratio $|\delta\omega_{0}/\omega_{0}|$, which weights it. The participation ratio is the ratio of energy stored in the sample component with respect to the total stored energy in the resonatorMinev et al. (2021). A high $Q_{0}$ is necessary to reliably extract the loss of the sample component from the measurement of the $Q_{i}$. Using Eqn. 6 and 10, we can rewrite Eqn. 9 as
@@ III.2 Half-wave resonator circuit model
i.e. $\delta\omega_{0}$ is positive (negative) if the element is inductive (capacitive), effectively shortening (elongating) the electrical length of the resonator—opposite the sign of the quarter-wave result (Eqn. 8).
@@
This $Q_{\rm eff}$ term has two contributions: the loss tangent of the sample component $\tan{\delta}$, and the participation ratio $|\delta\omega_{0}/\omega_{0}|$, which weights it. Moreover, the loss of the reactive element in a half-wave resonator $r$ effectively acts as a shunt resistor $R$ to that of resonator, i.e. $R_{0}$. Using Eqns. 12 and 15, we can rewrite Eqn. 16 as:
@@ IV.1 Kinetic inductance of Al strip
Fig. 2 displays the three coplanar waveguide resonators (from left to right): ‘reference’, ‘control’, and ‘sample’ resonators. The three high-$Q$ ($Q\sim 10^{5}$) resonators of prescribed lengths are capacitively coupled to a transmission feedline. The length of each resonator is designed to yield a resonance frequency in the range of 3–9 GHz, to match the cryogenic amplifier and circulator bandwidth. The choice of different resonance frequencies allows for measurement of multiple sample components per chip for each experimental cool-down. We also maximize the distance between reference and test resonator frequencies to help identify the correct resonator during measurement. When we are uncertain in identifying a resonance, we can temporarily eliminate the coplanar waveguide to the ground plane with a drop of silver epoxy, and recover it by dissolving the epoxy in an acetone bath.
The reference resonator on the far left in Fig. 2A is a half-wave resonator consisting only of Nb, with no additional material. Fig. 2B displays the capacitive coupling between the resonators and the transmission line, which sets the resonators’ coupling $Q$ factor $Q_{c}$. By measuring the frequency and $Q$ factor of the reference resonators as a function of temperature, we quantify any background kinetic inductance and dissipation from the Nb resonators. Panels C and D show the Al patterns of the sample component and control resonators, respectively. The control resonator is terminated directly to the ground plane without the long Al sample component that we intend to measure.
@@
To eliminate the effect of the microwave environment, including interference due to multiple reflections of the probe signal along the measurement chain, we analyze the $S_{21}$ data using the circle fitting methodProbst et al. (2015) to extract the resonance frequency as well as the internal quality factor $Q_{i}$ of the microwave resonator, which are plotted in Fig. 4A and B, respectively. When applying the equation for the frequency shift of the quarter-wave resonator in Table 1, we use the definition of $\Delta f=f(T)-f_{0}$ and $x=2\pi f_{0}L_{K}(T)$ where $f(T)$, $f_{0}$, and $L_{K}(T)$ are the resonance frequency at zero temperature, i.e. $T=0$ K, and the kinetic inductance of the Al, respectively. Using Eqn. 8, we have
@@
To fit Eqn. 18 to our data, we use BCS theory for the kinetic inductanceAnnunziata et al. (2010):
@@
where $\hbar$ is the reduced Planck’s constant, $k_{B}$ the Boltzmann constant, $T_{c}$ the critical temperature, $l=420~{}\mu$m the length, $w=10~{}\mu$m the width, $R_{\square}$ the sheet resistance, and $\Delta$ the BCS temperature-dependent gap energy of the Al strip. Using the approximation of $\Delta=\Delta_{0}[1-(T/T_{c})^{4}]^{1/2}$ with $\Delta_{0}=1.764\cdot k_{B}T_{c}$, we employ a least-squares fit of the data to Eqn. 18 with three parameters: $T_{c}$, $R_{s}$, and $f_{0}$, and plot the result as the brown line in Fig. 4A. We obtain $f_{0}$, the resonance frequency without the contribution of $L_{K}(T)$, from the fitting rather than the measured control resonator frequency because the resonance frequencies of lithographically identical resonators may vary by approximately 3 MHz, possibly due to slight fabrication differences or due to interaction between two resonances at the same frequency (see Fig. 10 in Appendix VII.3). This statistical variance is slightly less than one part per thousand (ppt), but sizable when compared to the shift of resonance frequency induced by the kinetic inductance. This uncertainty in $f_{0}$ constitutes the limitation in accuracy of this measurement method. However, this measurement method has high precision, inherited from the high-$Q$ factor of the resonators. As such, the error bar of the resonance frequency in Fig. 4A is less than one part per million (ppm).
Figure 4: (A) Resonance frequency (green data points) and (B) internal-$Q$ factor $Q_{i}$ of the quarter-wave resonator containing an Al sample component. (A) Solid brown curve is the best fit to BCS theory, i.e. Eqn. 18 and 19. Cooper pair density in Al decreases with increasing temperature, thus increasing kinetic inductance. Right y-axis portrays the data as kinetic inductance in units of pH and $\mathcal{D}_{s}$ in units of K. (B) Using the fitted parameters in (A), we calculate $Q_{i}$ according to the Mattis-Bardeen theory of surface impedance (solid brown line). The remarkable agreement to the data suggests the degradation of $Q_{i}$ for $T\geq 0.4$ K is dominated by thermally excited quasiparticles in the sample component rather than the resonator because the corresponding drop in $Q_{i}$ for the control resonator in the same temperature range is much smaller (blue dashed line). For $T<0.4$ K, however, $Q_{i}$ is potentially dominated by the resonator, not the Al strip.
@@
By measuring the quality factors of the bare resonator and ($Q_{0}$) and the sample-terminated resonator ($Q_{i}$), we can infer the quality factor of the sample component ($Q_{\rm{eff}}$). This value then yields the loss tangent $\delta$ (Eqn. 10) at the resonance frequency. The loss tangent $\tan{\delta}=\sigma_{1}/\sigma_{2}$ is equivalent to the ratio of the resistive and reactive components of the complex conductivity $\sigma=\sigma_{1}-j\sigma_{2}$. We note that here we follow the negative sign convention first introduced by Glover and TinkhamGlover and Tinkham (1957) wherein microwave photons were used to probe the complex conductivity. In fact, Glover and Tinkham reported observation of the superconducting energy gap in far-infrared measurements(submitted May 17, 1957) before the paper describing BCS theory, which formed the basis for the 1972 Nobel prize in physics, was submitted (July 8, 1957).
@@
Examining the data below 0.4 K carefully, we observe that the measured $Q_{\rm i}$ slighly exceeds that of the control resonator and shows a minor peak near 0.2 K. This peak in $Q_{i}$ is a key signature of the $Q$ factor being limited by dissipation through two-level systems (TLSs). At temperatures well below $\hbar\omega_{0}/k_{B}$, the TLSs are in their ground states and ready to absorb photon energy from the measurement source. However, at temperatures well above $\hbar\omega_{0}/k_{B}$, the TLSs are already thermally excited out of their ground states, therefore reducing the probability of photon absorption and subsequent losses through the TLS channel. As a result, $Q_{i}$ peaks at around $\hbar\omega_{0}/k_{B}$, i.e. 0.2 K for our sample component resonator, before its eventual degradation due to quasiparticles as temperature increasesAlexander et al. (2022). This minor peak in $Q_{i}$ is suppressed when measuring the resonator with higher power, because microwave photons from the measurement source can saturate the TLSs. The behavior of $Q_{i}$ as a function of temperature at various powers, plotted in Fig. 11 in Appendix VII.3, follows this expectation.
Figure 5: Optical image of the niobium diselenide sample component embedded in the quarter-wave resonator. (A) Grounding the coplanar waveguide resonator through an exfoliated strip of NbSe2 sample component using niobium nitride. The coplanar waveguide and the ground plane are made of Nb (light gray). (B) Close up image of the NbSe2 sample component in a light green color. The scale bars in panel (A) and (B) are 100 and 20 $\mu$m, respectively.

### v3

@@ I Introduction
Precise measurement of complex conductivity can act as a powerful tool to study the fundamental properties of superconductors Tinkham (1996). The reactive component of the conductivity is the kinetic inductance, which arises from the kinetic energy of Cooper pairs. Kinetic inductance can directly measure superfluid stiffness, and therefore, has the capability to reveal quantum geometric contributions to stiffness which are invisible at the level of band dispersion Tovmasyan et al. (2016); Julku et al. (2016); Verma et al. (2021); Törmä et al. (2022), and for which there only exists indirect evidence at present Tian et al. (2023) in twisted graphene. Sensitive measurements of the superfluid stiffness can also be used to detect the weak supercurrent dependence of the penetration depth (known as the nonlinear Meissner effect Yip and Sauls (1992); Xu et al. (1995); Stojkovi懠and Valls (1995); Sauls (2022)), which offers highly constraining insights into the superconducting pairing symmetry. On the other hand, the resistive component of the complex conductivity reflects the energy dissipation due to normal-state carriers under the framework of the two-fluid model. Finally, kinetic inductance is an important ingredient in superconducting quantum circuitry: it can be used to engineer qubits Winkel et al. (2020), to realize high-sensitivity photon detectors Day et al. (2003), and to parametrically amplify signals with performance constrained solely by quantum noise Eom et al. (2012); Ranzani et al. (2018).
@@
In this paper, we describe a new technique for accurately determining the kinetic inductance of these thin, small superconductor sample components. Inspired by the use of superconducting resonators in the readout of qubit states in circuit quantum electrodynamics (cQED), our method is a mesoscopic hybrid setup in which the kinetic inductance of our sample component acts as a small perturbation to the probe resonator. We then extract the relevant sample component parameters from the change of the resonator frequency and quality factor. This contrasts with previous kinetic inductance measurements in which a 50 $\Omega$ resistor shunts out the resonator in order to impedance match to the measurement circuitry, thus reducing the quality ($Q$) factor and sacrificing measurement sensitivity Annunziata et al. (2010); Singh et al. (2018). Our method shares the same fundamental mechanism as kinetic inductance detectors, which measure the suppression of resonance frequency due to the breaking of Cooper pairs in a superconducting resonator Visser et al. (2014); Day et al. (2003). Already, this method has been applied to study the superfluid stiffness in superconductor-ferromagnet bilayers BÞttcher et al. (2023), FeSe1-xTex Cherpak et al. (2023), and magic-angle twisted bilayer graphene Tanaka et al. (2024), as well as metrological studies of the loss tangent of materials and the cleanliness of fabrication procedures to advance the development of superconducting quantum technology Sage et al. (2011); Quintana et al. (2014); Grünhaupt et al. (2018); McRae et al. (2021); Antony et al. (2022).
@@ II Theory of Kinetic Inductance and Superfluid Stiffness
In a superconductor, kinetic inductance arises from the inertial response of Cooper pairs to an applied current Meservey and Tedrow (1969). If all the Cooper pairs contribute to the supercurrent, we can equate the total kinetic energy to kinetic inductance and derive
@@
Apart from the spatial profile of ${\bf A}({\bf r})$, we have assumed both the locality and isotropy of the Mattis-Bardeen kernel Mattis and Bardeen (1958) to relate the current density to the vector potential as explained in the Appendix VII.1:
$J_{\mu}({\bf r})\propto\mathcal{D}_{s}\delta_{\mu\nu}{\bf A}_{\nu}({\bf r}),$ (4)
where $\mu,\nu$ denote spatial directions. While the more general Mattis-Bardeen kernel offers a complete description, it complicates the analysis and obscures the relationship between kinetic inductance, penetration depth, and superfluid stiffness. Specifically, the superfluid stiffness is defined as the ${\bf q}=0$ component of the Fourier transform of the Mattis-Bardeen kernel, $K({\bf q})$ Tinkham (1996). The penetration depth is derived from the solution of the non-local London equation Tinkham (1996), whereas the kinetic inductance characterizes an effective penetration depth that depends on both $K({\bf q})$ and the boundary conditions Gao (2008). Next, we address each assumption individually.
As we discuss in Appendix VII.1, while the locality assumption might seem invalid for superconductors with large coherence lengths, practical factors such as sample thickness and electronic mean free path strongly reinforce its validity. One may heuristically think of an effective coherence length that is dominated by the smallest of the electronic mean free path and the sample thickness. As a result, despite the potential complications of a non-local kernel, the local approximation retains widespread applicability, particularly in thin samples.
In contrast, the breakdown of isotropic response in anisotropic superconductors is unavoidable. In these systems, the kinetic inductance depends on the alignment of current flow with the crystalline axes of the material. Beyond geometric factors related to the effective penetration volume (as detailed in Appendix VII.1), the kinetic inductance picks up the penetration depth along the direction of the current.
@@ III.1 Quarter-wave resonator circuit model
where $k$ is the wavenumber. The approximation in Eqn. 6 uses $|r+jx|\ll Z_{0}$, so that the sample component introduces only a small perturbation to the quarter-wave resonator impedance. By comparing the transmission line resonator with the parallel RLC resonator and relating their physical parameters (see Appendix VII.2), the loss becomes $\alpha\ell+r/Z_{0}$ with a frequency shift, $\delta\omega_{0}$, given by:
@@
The resonance-frequency shift is negative for an inductive sample, which effectively elongates the electrical length of the resonator.
@@
This $Q_{\rm eff}$ term has two contributions: the loss tangent of the sample component $\tan{\delta}$, and the participation ratio $|\delta\omega_{0}/\omega_{0}|$, which weights it. The participation ratio is the ratio of energy stored in the sample component with respect to the total stored energy in the resonator Minev et al. (2021). A high $Q_{0}$ is necessary to reliably extract the loss of the sample component from the measurement of the $Q_{i}$. Using Eqn. 5 and 9, we can rewrite Eqn. 8 as
@@ III.2 Half-wave resonator circuit model
i.e. $\delta\omega_{0}$ is positive (negative) if the element is inductive (capacitive), effectively shortening (elongating) the electrical length of the resonator—opposite the sign of the quarter-wave result (Eqn. 7).
@@
This $Q_{\rm eff}$ term has two contributions: the loss tangent of the sample component $\tan{\delta}$, and the participation ratio $|\delta\omega_{0}/\omega_{0}|$, which weights it. Moreover, the loss of the reactive element in a half-wave resonator $r$ effectively acts as a shunt resistor $R$ to that of resonator, i.e. $R_{0}$. Using Eqns. 11 and 14, we can rewrite Eqn. 15 as:
@@ IV.1 Kinetic inductance of Al strip
Fig. 2 displays the three coplanar waveguide resonators (from left to right): ‘reference’, ‘control’, and ‘sample’ resonators. The three high-$Q$ ($Q\sim 10^{5}$) resonators of prescribed lengths are capacitively coupled to a transmission feedline. The length of each resonator is designed to yield a resonance frequency in the range of 3–9 GHz, to match the cryogenic amplifier and circulator bandwidth. The choice of different resonance frequencies allows for measurement of multiple sample components per chip for each experimental cool-down. We also maximize the separation between reference and test resonator frequencies to help identify the correct resonator during measurement. When we are uncertain in identifying a resonance, we can temporarily eliminate the coplanar waveguide to the ground plane with a drop of silver epoxy, and recover it by dissolving the epoxy in an acetone bath.
The reference resonator on the far left in Fig. 2A is a half-wave resonator consisting only of Nb, with no additional material. Fig. 2B displays the capacitive coupling between the resonators and the transmission line, which sets the resonators’ coupling $Q$ factor $Q_{c}$. By measuring the resonance frequency and $Q$ factor of the reference resonators as a function of temperature, we quantify any background kinetic inductance and dissipation from the Nb resonators. Panels C and D show the Al patterns of the sample component and control resonators, respectively. The control resonator is terminated directly to the ground plane without the long Al sample component that we intend to measure. The role of the control resonator is to monitor the temperature dependence of the frequency and quality factor of the niobium resonator used to measure our sample.
@@
To eliminate the effect of the microwave environment, including interference due to multiple reflections of the probe signal along the measurement chain, we analyze the $S_{21}$ data using the circle fitting method Probst et al. (2015) to extract the resonance frequency as well as the internal quality factor $Q_{i}$ of the microwave resonator, which are plotted in Fig. 4A and B, respectively. When applying the equation for the frequency shift of the quarter-wave resonator in Table 1, we use the definition of $\Delta f=f(T)-f_{0}$ and $x=2\pi f_{0}L_{K}(T)$ where $f(T)$, $f_{0}$, and $L_{K}(T)$ are the resonance frequency at zero temperature, i.e. $T=0$ K, and the kinetic inductance of the Al, respectively. Using Eqn. 7, we have
@@
To fit Eqn. 17 to our data, we use BCS theory for the kinetic inductance Annunziata et al. (2010):
@@
where $\hbar$ is the reduced Planck’s constant, $k_{B}$ the Boltzmann constant, $T_{c}$ the critical temperature, $l=420~{}\mu$m the length, $w=10~{}\mu$m the width, $R_{\square}$ the sheet resistance, and $\Delta$ the BCS temperature-dependent gap energy of the Al strip. Using the approximation of $\Delta=\Delta_{0}[1-(T/T_{c})^{4}]^{1/2}$ with $\Delta_{0}=1.764\cdot k_{B}T_{c}$, we employ a least-squares fit of the data to Eqn. 17 with three parameters: $T_{c}$, $R_{s}$, and $f_{0}$, and plot the result as the brown line in Fig. 4A. We obtain $f_{0}$, the resonance frequency without the contribution of $L_{K}(T)$, from the fitting rather than the measured control resonator frequency because the resonance frequencies of lithographically identical resonators may vary by approximately 3 MHz, possibly due to slight fabrication differences or due to interaction between two resonances at the same frequency (see Fig. 10 in Appendix VII.3). This statistical variance is slightly less than one part per thousand (ppt), but sizable when compared to the shift of resonance frequency induced by the kinetic inductance. This uncertainty in $f_{0}$ constitutes the limitation in accuracy of this measurement method. However, this measurement method has high precision, inherited from the high-$Q$ factor of the resonators. As such, the error bar of the resonance frequency in Fig. 4A is less than one part per million (ppm).
Figure 4: (A) Resonance frequency (green data points) and (B) internal-$Q$ factor $Q_{i}$ of the quarter-wave resonator containing an Al sample component. (A) Solid brown curve is the best fit to BCS theory, i.e. Eqn. 17 and 18. Cooper pair density in Al decreases with increasing temperature, thus increasing kinetic inductance. Right y-axis portrays the data as kinetic inductance in units of pH and $\mathcal{D}_{s}$ in units of K. (B) Using the fitted parameters in (A), we calculate $Q_{i}$ according to the Mattis-Bardeen theory of surface impedance (solid brown line). The remarkable agreement to the data suggests the degradation of $Q_{i}$ for $T\geq 0.4$ K is dominated by thermally excited quasiparticles in the sample component rather than the resonator because the corresponding drop in $Q_{i}$ for the control resonator in the same temperature range is much smaller (blue dashed line). For $T<0.4$ K, however, $Q_{i}$ is potentially dominated by the resonator, not the Al strip.
@@
By measuring the quality factors of the bare resonator and ($Q_{0}$) and the sample-terminated resonator ($Q_{i}$), we can infer the quality factor of the sample component ($Q_{\rm{eff}}$). This value then yields the loss tangent $\delta$ (Eqn. 9) at the resonance frequency. The loss tangent $\tan{\delta}=\sigma_{1}/\sigma_{2}$ is equivalent to the ratio of the resistive and reactive components of the complex conductivity $\sigma=\sigma_{1}-j\sigma_{2}$. We note that here we follow the negative sign convention first introduced by Glover and Tinkham Glover and Tinkham (1957) wherein microwave photons were used to probe the complex conductivity. In fact, Glover and Tinkham reported observation of the superconducting energy gap in far-infrared measurements(submitted May 17, 1957) before the paper describing BCS theory, which formed the basis for the 1972 Nobel prize in physics, was submitted (July 8, 1957).
@@
Examining the data below 0.4 K carefully, we observe that the measured $Q_{\rm i}$ slightly exceeds that of the control resonator and shows a minor peak near 0.2 K. This peak in $Q_{i}$ is a key signature of the $Q$ factor being limited by dissipation through two-level systems (TLSs). At temperatures well below $\hbar\omega_{0}/k_{B}$, the TLSs are in their ground states and ready to absorb photon energy from the measurement source. However, at temperatures well above $\hbar\omega_{0}/k_{B}$, the TLSs are already thermally excited out of their ground states, therefore reducing the probability of photon absorption and subsequent losses through the TLS channel. As a result, $Q_{i}$ peaks at around $\hbar\omega_{0}/k_{B}$, i.e. 0.2 K for our sample component resonator, before its eventual degradation due to quasiparticles as temperature increases Alexander et al. (2022). This minor peak in $Q_{i}$ is suppressed when measuring the resonator with higher power, because microwave photons from the measurement source can saturate the TLSs. The behavior of $Q_{i}$ as a function of temperature at various powers, plotted in Fig. 11 in Appendix VII.3, follows this expectation.
Figure 5: Optical image of the niobium diselenide sample embedded in the quarter-wave resonator. (A) Grounding the coplanar waveguide resonator through an exfoliated strip of NbSe2 sample using niobium nitride. The coplanar waveguide and the ground plane are made of Nb (light gray). (B) Close up image of the NbSe2 sample in a light green color. The scale bars in panel (A) and (B) are 100 and 20 $\mu$m, respectively.